OpenStudy (anonymous):

Solve the following quadratic equation by a method of your choice: y2 – 16 = 0. A. y = ±4 B. y = ±16 C. y = -2 or 3 D. y = 1 or 5

OpenStudy (anonymous):

y^2 = 16 sqrt(y^2) = sqrt(16) y = +/- 4 A

OpenStudy (anonymous):

It is right Go to www.wolframalpha.com